OpenStudy (anonymous):

3 over 4 - 7 over 10

OpenStudy (anonymous):

well first, you will want to give them the same denominator so you can do 15/20 - 14/20

OpenStudy (anonymous):

Take the denomantors times 20 because both 4 and 10 can go into 20 then take the nemorators and multiple them by 20 then divide so 20 * 3 = 60 then divided by 4 minus 20 * 7 = 140 divided by 10 so 15- 14 = 1